The Large-Scale Industrial Corporation, a large well established company, is about to issue $6,000,000 worth of commercial paper. The paper has a maturity of 9 months (270 days), and commands a price worth 97.5% of par value in the market. The paper will be sold with a face or par value of $100,000.

How many commercial papers will be sold?

What is the cost of this borrowing to the firm?

please answer the problem with the steps and formula....thank you
Answer #1

Proceeds from the issue of the commercial paper = $6,000,000 * 97.5% = 5,850,000

Cost of borrowing over 270 days = ( 6,000,000 - 5,850,000) / 5,850,000* 365/ 270 = 0.0347 or 3.47%

Effective annual rate of borrowing = (1+ 3.47%) (365/270) -1 = 3.48%

No of commercial papers = 6,000,000 / 100,000= 60 papers